OpenStudy (anonymous):

is 17 irrational? just a quick question and can someone help me?

OpenStudy (igreen):

Rational numbers can be written as simple fractions.

OpenStudy (igreen):

Can you write 17 as a simple fraction?

OpenStudy (yanasidlinskiy):

What's the difference between rational and irrational?

OpenStudy (yanasidlinskiy):

Ahh, iGreen, lololol.

OpenStudy (igreen):

Haha, too fast for you :P

OpenStudy (yanasidlinskiy):

:P

OpenStudy (anonymous):

17 is irrational because nothing goes into it

OpenStudy (igreen):

That is not the definition of an irrational number..

OpenStudy (rational):

you're confusing primes with irrationals

OpenStudy (igreen):

That would be saying it's prime

OpenStudy (anonymous):

17 is a irrational number

OpenStudy (anonymous):

ohh

OpenStudy (rational):

\[\large 17 ~~= ~~\dfrac{17}{1}~~\color{blue}{\text{ratio of two integers}}\] so \(17\) is rational

OpenStudy (anonymous):

so 17 is rational?

OpenStudy (rational):

that also means \(17\) is NOT irrational

OpenStudy (anonymous):

@iGreen is he correct?

OpenStudy (anonymous):

yes 17 is rational because it can be ritten in the form of p\q where p and q are integers and q is not equal to 0

OpenStudy (anonymous):

written*

OpenStudy (igreen):

Yes..he is 100% correct @pjpineda17

OpenStudy (anonymous):

theses are my options Which number is irrational? A. 1 over 6 B. square root of 17 C. square root of 49 D. Syntax error from line 1 column 53 to line 1 column 60. Unexpected '0'.

OpenStudy (anonymous):

B

OpenStudy (anonymous):

D is 0.06 repeated

OpenStudy (anonymous):

\[\sqrt{17}\]

OpenStudy (anonymous):

is irrational =4.123105626

OpenStudy (anonymous):

and it is non terminating

OpenStudy (anonymous):

so its B

OpenStudy (rational):

Yes \(\sqrt{17}\) is irrational